# Universal properties of the evolution of the Universe in modified loop quantum cosmology

### v2

@@ Abstract
In this paper, we systematically study the evolution of the Universe in the framework of a modified loop quantum cosmological model (mLQC-I) with various inflationary potentials, including chaotic, Starobinsky, generalized Starobinsky, polynomials of the first and second kinds, generalized T-models and natural inflation. In all these models, the big bang singularity is represented by a quantum bounce, and the evolution of the Universe both before and after the bounce is universal and weakly depends on the inflationary potentials, as long as the evolution is dominated by the kinetic energy of the inflaton at the bounce. In particular, the evolution in the pre-bounce region can be universally divided into three different phases: pre-bouncing, pre-transition, and pre-de Sitter. The pre-bouncing phase occurs immediately before the quantum bounce, during which the evolution of the Universe is dominated by the kinetic energy of the inflaton. Thus, the equation of state of the inflaton is about one, $w(\phi)\simeq 1$. Soon, the inflation potential takes over, so $w(\phi)$ rapidly falls from one to negative one. This pre-transition phase is very short and quickly turns into the pre-de Sitter phase, whereby the effective cosmological constant with a Planck size takes over and dominates the rest of the contracting phase. In the entire pre-bounce regime, the evolution of the expansion factor and the inflaton can be approximated by analytical solutions, which are universal and independent of the inflation potentials.
@@ I Introduction
Then, a natural question arises: What happened before inflation? This important question is expected to be answered by quantum gravity (QG), a theory that has not yet been established, despite enormous efforts in the past century, although many candidates have been proposed. Among them are string/M-theory Green et al. (2012); Becker et al. (2006) and loop quantum gravity (LQG) Ashtekar and Lewandowski (2004); Thiemann (2007); Bojowald (2010); Gambini and Pullin (2011); Rovelli and Vidotto (2014a). In particular, in the last two decades, LQG has been rigorously applied to understand singularity resolution in various cosmological models (for recent reviews, see Refs. Ashtekar and Singh (2011); Elizaga Navascués and Marugán (2021); Li and Singh (2023); Agulló et al. (2023)). In all these models, a coherent picture of Planck scale physics has emerged: the big bang singularity is replaced by a quantum bounce, purely due to quantum geometric effects. This framework is often referred to as loop quantum cosmology (LQC).
@@
The physical implications of LQC have also been studied using the effective descriptions of the quantum spacetime derived from coherent states Taveras (2008), whose validity has been numerically verified for various spacetime Singh (2018), especially for states that sharply peaked on classical trajectories at late times Kamiński et al. (2020). Effective dynamics provide a definitive answer to the resolution of the big bang singularity, replaced by a quantum bounce when the energy density of matter reaches a maximum value determined purely by the underlying quantum geometry Ashtekar and Singh (2011); Elizaga Navascués and Marugán (2021); Li and Singh (2023); Agulló et al. (2023).
@@
Despite a wealth of results on the singularity resolution and phenomenology of the very early universe obtained in the framework of LQC, an important issue that has remained open is its connection to LQG (see, for example, Beetle et al. (2017); Bojowald (2021) for discussions). The starting point of LQC is to first classically reduce the Hamiltonian from infinitely many to a few gravitational degrees of freedom by imposing homogeneity and then to quantize the classically reduced Hamiltonian using the techniques of LQG. However, in LQG, the processes of symmetry reduction and quantization do not commute in general, and it is important to understand how well the physics of the full LQG is captured by LQC. In addition, in LQG, the Hamiltonian usually consists of two parts, the Euclidean and Lorentz parts, and one follows different processes to quantize each part Ashtekar and Lewandowski (2004); Thiemann (2007); Bojowald (2010); Gambini and Pullin (2011); Rovelli and Vidotto (2014a). However, in LQC, only the quantization of the Euclidean part was considered by taking advantage of the properties of the classical Hamiltonian, in which the Euclidean and Lorentz parts are proportional to each other for the flat Friedmann-Lemaítre-Robertson-Walker (FLRW) universe Ashtekar and Singh (2011).
@@
In this paper, we shall systematically study the evolution of the Universe in the framework of mLQC-I. Note that such studies have already been conducted in the post-bounce regime ($t\geq t_{B}$) in Li et al. (2018a, b, 2019a, 2021), where $t_{B}$ denotes the bounce time. It was found that the evolution in this regime is universal and independent of the inflationary potential, as long as the evolution is dominated at the quantum bounce by the kinetic energy of the inflaton
@@
where an over dot denotes the derivative with respect to the cosmic time, and $\phi_{B}$ is the value of the inflaton at the bounce. Initial conditions satisfying Eq.(1.1) are important because they always lead to slow-roll inflation Li et al. (2018a, b, 2019a, 2021), quick similar to that in LQC Bonga and Gupt (2016); Zhu et al. (2017). Therefore, in this paper, we focus on the pre-bounce regime ($t\leq t_{B}$).
Specifically, the rest of the paper is organized as follows: In Sec.II we briefly review the mLQC-I model and write down the dynamical Hamiltonian equations. To understand the major properties of mLQC-I, we also write down the corresponding FR equations. In Sec.III, we study the evolution of the Universe with several inflationary potentials, including chaotic, Starobinsky, generalized Starobinsky, polynomials of the first and second kinds, generalized T-models and natural inflation. In the last case, we consider several choices of parameters involved in the model. Note that the chaotic and natural inflationary models have some tensions with current observations Akrami et al. (2020). The reasons that we still consider them here are two-folds: First, after quantum gravitational effects are taken into account, such tensions can be alleviated Zhu et al. (2016). Second, the universal properties of the evolution of the universe both before and after the quantum bounce are independent of the choice of inflationary potentials in LQC Zhu et al. (2017); Shahalam et al. (2017, 2018); Sharma et al. (2018, 2019); Shahalam et al. (2020); Levy and O. Ramos (2024). We find that this is also true in mLQC-I. In all these models, we find that the evolution of the Universe in the pre-bounce regime ($t\leq t_{B}$) is universal and weakly depends on the inflationary potentials, as long as the kinetic-energy-dominated condition (1.1) is satisfied at the bounce. In all these models, the evolution is clearly divided into three different epochs, pre-bouncing, pre-transition, and pre-de Sitter, as defined by Eq(II), in terms of the equation of state of the inflaton
@@
where $P(\phi)$ and $\rho(\phi)$ are the pressure and energy density of the inflationary field, respectively, given by Eq.(II). Moreover, in these epochs the expansion factor $a(t)$ and the inflaton $\phi(t)$ are universal and weakly depend on the initial conditions $\phi_{B}$ as well as on the inflationary potentials as shown explicitly by Fig. 31 of Sec. IV. In this section, we also show that $a(t)$ and $\phi(t)$ can be well-approximated by Eqs.(3.16) and (3.19), respectively, where $d_{n}$ and $e_{n}$ are constants determined by fitting the analytical expressions with their numerical solutions, as given in Sec. IV. Therefore, the evolution of the Universe in the pre-bounce regime ($t\leq t_{B}$) is also universal, quite similar to that in the post-bounce regime ($t\geq t_{B}$) Li et al. (2018a, b, 2019a, 2021). This is important and will significantly facilitate the studies of cosmological perturbations Li et al. (2020a, b, 2024), as already shown in LQC Zhu et al. (2017). The paper ends in Sec.V, in which our main results and concluding remarks are presented.
@@ II Effective Dynamics of Modified Loop Quantum Cosmology
where $\gamma$ is known as the Barbero-Immirzi parameter whose value is set to $\gamma\approx 0.2375$ using black hole thermodynamics in LQG Meissner (2004). The parameter $\lambda$ is defined as $\lambda^{2}\equiv\Delta=4\sqrt{3}\pi\gamma\ell^{2}_{Pl}$, where $\Delta$ denotes the minimal area of the area operator in LQG Ashtekar and Lewandowski (2004); Thiemann (2007); Bojowald (2010); Gambini and Pullin (2011); Rovelli and Vidotto (2014b); Ashtekar and Pullin (2017).
@@
Figure 1: The evolution $w(t)$ of the equation of state in the post-bounce phase ($t\geq t_{B}$) for the chaotic potential with different initial conditions $\phi_{B}$ for $\dot{\phi}_{B}>0$. In this plot, $t_{B}$ is set to zero.
It should be noted that in Fig. 1, only the case of the chaotic potential was considered. However, the general behavior of the evolution of the universe is universal and independent of the choice of potential, as long as the initial conditions are kinetic energy dominated. These remarkable features are also shared by LQC, first found in Zhu et al. (2017) and later confirmed in various cases Shahalam et al. (2017, 2018); Sharma et al. (2018, 2019); Shahalam et al. (2020).
@@ III Numerical solutions of the Evolution of the Universe
    To have a successful slow-roll inflation, the inflation potential has to be very flat, so that the Universe can expand large enough Baumann (2011). All the cosmological problems can be resolved if the Universe expands about 60 e-fold during the inflationary phase, although its exact value depends on the inflationary models.
@@ III.3 $\alpha$- attractor Potentials
Figure 11: The equation of state for the generalized Starobinsky potentials (3.11) for different values of $\phi_{B}$ that result in at least 50 e-Folds with $\alpha=0.0962$ and $V_{0}=10^{-12}$ GeV. (a) for $\dot{\phi}_{B}>0$ and (b) for $\dot{\phi}_{B}<0$.
@@
Figure 12: The numerical solutions of $a(t)$ and $\phi(t)$ for the $\alpha$-attractor potential (3.11) with $\alpha=0.0962$ and $V_{0}=10^{-12}$ GeV. (a) For $\dot{\phi}_{B}>0$, and (b) for $\dot{\phi}_{B}<0$.
@@
Figure 13: E-fold $N$ for different values of $\phi_{B}$ for the polynomial potential of the first kind given by Eq.(3.12) where $V_{0}=3.3787*10^{-15}$ GeV and $\mu=0.31075$. Left Panel: $\dot{\phi}_{B}>0$. Right Panel: $\dot{\phi}_{B}<0$.
@@
Figure 14: The equation of state $w(t)$ with different values of $\phi_{B}$ for for the polynomial potential of the first kind given by Eq.(3.12) with $V_{0}=3.3787\times 10^{-15}$ GeV and $\mu=0.31075$. (a) $\dot{\phi}_{B}>0$ and (b) $\dot{\phi}_{B}<0$.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 15: The numerical solutions of $a(t)$ and $\phi(t)$ with different initial values of $\phi_{B}$ for the polynomial potential of the first kind given by Eq.(3.12), where $V_{0}=3.3787\times 10^{-15}$ GeV and $\mu=0.31075$. (a) For $\dot{\phi}_{B}>0$ and (b) for $\dot{\phi}_{B}<0$.
Refer to caption Refer to caption
Figure 16: E-fold $N$ for different values of $\phi_{B}$ for the polynomial potential of the second kind given by Eq.(3.13) with $V_{0}=3.2599\times 10^{-15}$ GeV and $\mu=0.01043$. Left Panel: $\dot{\phi}_{B}>0$. Right Panel: $\dot{\phi}_{B}<0$.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 17: The equation of state $w(t)$ with different values of $\phi_{B}$ for for the polynomial potential of the second kind given by Eq.(3.13) with $V_{0}=3.2599\times 10^{-15}$ GeV and $\mu=0.01043$, where (a) $\dot{\phi}_{B}>0$ and (b) $\dot{\phi}_{B}<0$.
Refer to caption Refer to caption
@@
Figure 19: E-folds $N$ for the generalized T-Models given by (3.14) with $V_{0}=1.96\times 10^{-15}$ and $\alpha=0.0962$. Left Panel: $\dot{\phi}_{B}>0$. Right Panel: $\dot{\phi}_{B}<0$.
@@
Refer to caption Refer to caption
Figure 20: The equation of state $w(t)$ with different values of $\phi_{B}$ for for the the generalized T-Models given by (3.14) with $V_{0}=1.96\times 10^{-15}$ and $\alpha=0.0962$, where (a) $\dot{\phi}_{B}>0$ and (b) $\dot{\phi}_{B}<0$.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 21: The numerical solutions of $a(t)$ and $\phi(t)$ for various initial conditions of $\phi_{B}$ for the generalized T-Models given by (3.14) with $V_{0}=1.96*10^{-15}$ and $\alpha=0.0962$. (a) For with $\dot{\phi}_{B}>0$ and (b) for $\dot{\phi}_{B}<0$.

### v3

@@ Abstract
In this paper, we systematically study the evolution of the Universe in the framework of a modified loop quantum cosmological model (mLQC-I) with various inflationary potentials, including chaotic, Starobinsky, generalized Starobinsky, polynomials of the first and second kinds, generalized T-models and natural inflation. In all these models, the big bang singularity is replaced by a quantum bounce, and the evolution of the Universe both before and after the bounce is universal and weakly depends on the inflationary potentials, as long as the evolution is dominated by the kinetic energy of the inflaton at the bounce. In particular, the evolution in the pre-bounce region can be universally divided into three different phases: pre-bouncing, pre-transition, and pre-de Sitter. The pre-bouncing phase occurs immediately before the quantum bounce, during which the evolution of the Universe is dominated by the kinetic energy of the inflaton. Thus, the equation of state of the inflaton is about one, $w(\phi)\simeq 1$. Soon, the inflation potential takes over, so $w(\phi)$ rapidly falls from one to negative one. This pre-transition phase is very short and quickly turns into the pre-de Sitter phase, whereby the effective cosmological constant with a Planck size takes over and dominates the rest of the contracting phase. In the entire pre-bounce regime, the evolution of the expansion factor and the inflaton can be approximated by analytical solutions, which are universal and independent of the inflation potentials.
@@ I Introduction
Then, a natural question arises: What happened before inflation? This important question is expected to be answered by quantum gravity (QG), a theory that has not yet been established, despite enormous efforts in the past century, although many candidates have been proposed. Among them are string/M-theory Green et al. (2012); Becker et al. (2006) and loop quantum gravity (LQG) Ashtekar and Lewandowski (2004); Thiemann (2007); Bojowald (2010); Gambini and Pullin (2011); Rovelli and Vidotto (2014). In particular, in the last two decades, LQG has been rigorously applied to understand singularity resolution in various cosmological models (For reviews, see Refs. Bojowald (2005); Ashtekar and Singh (2011); Ashtekar and Barrau (2015); Agullo and Singh (2017); Wilson-Ewing (2017); Elizaga Navascués and Marugán (2021); Ashtekar and Bianchi (2021); Li and Singh (2023); Agulló et al. (2023)). In all these models, a coherent picture of Planck scale physics has emerged: the big bang singularity is replaced by a quantum bounce  [^2] Bouncing universe inspired by string theory was first proposed in Ref. Gasperini and Veneziano (1993), and later studied intensively by various authors, see, for example Refs. Gasperini et al. (1997); Gasperini and Veneziano (2003); Haro (2013); Alonso-Serrano et al. (2023); Conzinu et al. (2023) and references therein. Other bouncing models have been also studied extensively Khoury et al. (2001); Brown et al. (2008); Battefeld and Peter (2015); Brandenberger and Peter (2017); Ijjas and Steinhardt (2018); Chandran and Shankaranarayanan (2024). In this paper, we shall focus ourselves on the quantum bounce from LQG., purely due to quantum geometric effects. This framework is often referred to as loop quantum cosmology (LQC).
@@
The physical implications of LQC have also been studied using the effective descriptions of the quantum spacetime derived from coherent states Taveras (2008), whose validity has been verified for various spacetimes both numerically Singh (2018) and Corichi and Singh (2008), especially for states that sharply peaked on classical trajectories at late times Kamiński et al. (2020). Effective dynamics provide a definitive answer to the resolution of the big bang singularity, replaced by a quantum bounce when the energy density of matter reaches a maximum value determined purely by the underlying quantum geometry Ashtekar and Singh (2011); Elizaga Navascués and Marugán (2021); Li and Singh (2023); Agulló et al. (2023).
@@
Despite a wealth of results on the singularity resolution and phenomenology of the very early universe obtained in the framework of LQC, an important issue that has remained open is its connection to LQG (see, for example, Beetle et al. (2017); Bojowald (2021) for discussions). The starting point of LQC is to first classically reduce the Hamiltonian from infinitely many to a few gravitational degrees of freedom by imposing homogeneity and then to quantize the classically reduced Hamiltonian using the techniques of LQG. However, in LQG, the processes of symmetry reduction and quantization do not commute in general, and it is important to understand how well the physics of the full LQG is captured by LQC. In addition, in LQG, the Hamiltonian usually consists of two parts, the Euclidean and Lorentz parts, and one follows different processes to quantize each part Ashtekar and Lewandowski (2004); Thiemann (2007); Bojowald (2010); Gambini and Pullin (2011); Rovelli and Vidotto (2014). However, in LQC, only the quantization of the Euclidean part was considered by taking advantage of the properties of the classical Hamiltonian, in which the Euclidean and Lorentz parts are proportional to each other for the flat Friedmann-Lemaítre-Robertson-Walker (FLRW) universe Ashtekar and Singh (2011).
@@
In this paper, we shall systematically study the evolution of the Universe in the framework of mLQC-I. Note that such studies have already been conducted in the post-bounce regime ($t\geq t_{B}$) in Li et al. (2018a, b); de Haro (2018); Saini and Singh (2019a); Li et al. (2019a); Saini and Singh (2019b); Li et al. (2021); Li and Singh (2022a), where $t_{B}$ denotes the bounce time. It was found that the evolution in this regime is universal and independent of the inflationary potential, as long as the evolution is dominated at the quantum bounce by the kinetic energy of the inflaton
@@
where an over dot denotes the derivative with respect to the cosmic time, and $\phi_{B}$ is the value of the inflaton at the bounce. Initial conditions satisfying Eq.(1.1) are important because they always lead to slow-roll inflation Li et al. (2018a, b, 2019a, 2021), quick similar to that in LQC Bonga and Gupt (2016); Zhu et al. (2017a, b); Shahalam et al. (2017, 2018); Sharma et al. (2018, 2019); Shahalam et al. (2020). Therefore, in this paper, we focus on the pre-bounce regime ($t\leq t_{B}$), which has not been studied adequately so far and is important, especially when the initial conditions of cosmological perturbations are imposed in the contracting phase Ashtekar and Singh (2011); Elizaga Navascués and Marugán (2021); Li and Singh (2023); Agulló et al. (2023).
Specifically, the rest of the paper is organized as follows: In Sec.II we briefly review the mLQC-I model and write down the dynamical Hamiltonian equations. To understand the major properties of mLQC-I, we also write down the corresponding FR equations. In Sec.III, we study the evolution of the Universe with several inflationary potentials, including chaotic, Starobinsky, generalized Starobinsky, polynomials of the first and second kinds, generalized T-models and natural inflation. In the last case, we consider several choices of parameters involved in the model. Note that the chaotic and natural inflationary models have some tensions with current observations Akrami et al. (2020). The reasons that we still consider them here are two-folds: First, after quantum gravitational effects are taken into account, such tensions can be alleviated Zhu et al. (2016). Second, the universal properties of the evolution of the universe both before and after the quantum bounce are independent of the choice of inflationary potentials in LQC Zhu et al. (2017a, b); Shahalam et al. (2017, 2018); Sharma et al. (2018, 2019); Shahalam et al. (2020); Levy and O. Ramos (2024). We find that this is also true in mLQC-I. In all these models, we find that the evolution of the Universe in the pre-bounce regime ($t\leq t_{B}$) is universal and weakly depends on the inflationary potentials, as long as the kinetic-energy-dominated condition (1.1) is satisfied at the bounce. In all these models, the evolution is clearly divided into three different epochs, pre-bouncing, pre-transition, and pre-de Sitter, as defined by Eq.(III.1), in terms of the equation of state of the inflaton
@@
where $P(\phi)$ and $\rho(\phi)$ are the pressure and energy density of the inflationary field, respectively, given by Eq.(II). Physically, it can be understood as follows: In the pre-bouncing phase, the evolution of the universe is dominated by the kinetic energy as determined by the initial conditions imposed at the bounce, so it is expected that its evolution in this phase is independent of the inflationary potentials and $w(\phi)\simeq+1$. However, an effective Planck-size cosmological constant exists in the pre-bounce regime [cf. Eq.(2.19)], so it will soon dominate the evolution once apart from the bounce, during which we have $w(\phi)\simeq-1$. Clearly, during this phase the evolution of the universe will be also independent of the inflationary potentials. In between these two phases, a short pre-transition phase exists, during which $w(\phi)$ drops rapidly from $+1$ to $-1$ and behaves like a step function [^3] It should be noted that such universal properties in the pre-bounce regime may not be shared by the standard LQC Bonga and Gupt (2016); Zhu et al. (2017a, b); Shahalam et al. (2017, 2018); Sharma et al. (2018, 2019); Shahalam et al. (2020), as in LQC such an effective Planck-size cosmological constant is absent in the pre-bounce regime Ashtekar and Singh (2011); Elizaga Navascués and Marugán (2021); Li and Singh (2023); Agulló et al. (2023). As a result, in these epochs the expansion factor $a(t)$ and the inflaton $\phi(t)$ are universal, as shown explicitly by Fig. 28 of Sec. IV. In this section, we also show that $a(t)$ and $\phi(t)$ can be well-approximated by Eqs.(3.18) and (3.21), respectively, where $d_{n}$ and $e_{n}$ are constants determined by fitting the analytical expressions with their numerical solutions, as given in Sec. IV. Therefore, the evolution of the Universe in both of the pre-bounce regime ($t\leq t_{B}$) and the post-bounce regime ($t\geq t_{B}$) is universal in mLQC-I Li et al. (2018a, b, 2019a, 2021). This is important and will significantly facilitate the studies of cosmological perturbations Li et al. (2020a, b, 2024); Elizaga Navascués and Marugán (2021); Li and Singh (2022b), as already shown in LQC Zhu et al. (2017b). The paper ends in Sec.V, in which our main results and concluding remarks are presented.
@@ II Effective Dynamics of Modified Loop Quantum Cosmology
where $\gamma$ is known as the Barbero-Immirzi parameter whose value is set to $\gamma\approx 0.2375$ using black hole thermodynamics in LQG Meissner (2004). The parameter $\lambda$ is defined as $\lambda^{2}\equiv\Delta=4\sqrt{3}\pi\gamma\ell^{2}_{Pl}$, where $\Delta$ denotes the minimal area gap of the area operator in LQG Rovelli and Smolin (1995); Ashtekar and Lewandowski (1997); Thiemann (2007); Ashtekar and Singh (2011).
@@
Figure 1: The evolution $w(t)$ of the equation of state in the post-bounce phase ($t\geq t_{B}$) for the chaotic potential with different initial conditions $\phi_{B}$ for $\dot{\phi}_{B}>0$. In this plot, $t_{B}$ is set to zero. It should be noted that the equation of state, $w(t)$, behaves the same not only for $\dot{\phi}_{B}<0$ with the same chaotic potential, but also for other potentials, as long as the evolution of the universe at the bounce is kinetic-energy dominated Li et al. (2018a, b, 2019a, 2021).
It should be noted that in Fig. 1, only the case of the chaotic potential was considered. However, the general behavior of the evolution of the universe is universal and independent of the choice of potential, as long as the initial conditions are kinetic energy dominated. These remarkable features are also shared by LQC, first found in Zhu et al. (2017b) and later confirmed in various cases Shahalam et al. (2017, 2018); Sharma et al. (2018, 2019); Shahalam et al. (2020).
@@ III Numerical solutions of the Evolution of the Universe
    To have a successful slow-roll inflation, the inflation potential has to be very flat, so that the Universe can expand large enough Baumann (2011). All the cosmological problems can be resolved if the Universe expands about 60 e-fold during the inflationary phase, although its exact value depends on the inflationary models Akrami et al. (2020). Therefore, in the following one will see that the minimal e-fold $N_{\text{min}}$ will be different in different models.
@@ III.3 $\alpha$- attractor Potentials
Figure 11: The equation of state for the generalized Starobinsky potentials (3.11) for different values of $\phi_{B}$ that result in at least 50 e-Folds with $\alpha=0.0962$ and $V_{0}=10^{-12}$ GeV, where (a) for $\dot{\phi}_{B}>0$ and (b) for $\dot{\phi}_{B}<0$.
@@
Figure 12: The numerical solutions of $a(t)$ and $\phi(t)$ for the generalized Starobinsky potentials (3.11) with $\alpha=0.0962$ and $V_{0}=10^{-12}$ GeV. (a) For $\dot{\phi}_{B}>0$, and (b) for $\dot{\phi}_{B}<0$.
Figure 13: E-fold $N$ for $\dot{\phi}_{B}>0$ with different values of $\phi_{B}$ for the polynomial potential of the first kind given by Eq.(3.12) where $V_{0}=3.3787*10^{-15}$ GeV and $\mu=0.31075$. The case with $\dot{\phi}_{B}<0$ can be obtained by the symmetry (3.7).
@@
Figure 14: The equation of state $w(t)$ for the polynomial potential of the first kind given by Eq.(3.12) with $\dot{\phi}_{B}>0$ and different values of $\phi_{B}$, where $V_{0}=3.3787\times 10^{-15}$ GeV and $\mu=0.31075$.
@@
Figure 15: The numerical solutions of $a(t)$ and $\phi(t)$ with different initial values of $\phi_{B}$ for the polynomial potential of the first kind given by Eq.(3.12) with $\dot{\phi}_{B}>0$, where $V_{0}=3.3787\times 10^{-15}$ GeV and $\mu=0.31075$.
Figure 16: E-fold $N$ for different values of $\phi_{B}$ for the polynomial potential of the second kind given by Eq.(3.14) with $\dot{\phi}_{B}>0$, $V_{0}=3.2599\times 10^{-15}$ GeV and $\mu=0.01043$.
@@
Figure 17: The equation of state $w(t)$ with different values of $\phi_{B}$ for for the polynomial potential of the second kind given by Eq.(3.14) with $\dot{\phi}_{B}>0$, $V_{0}=3.2599\times 10^{-15}$ GeV and $\mu=0.01043$.
@@
Figure 19: E-folds $N$ for the generalized T-Models given by (3.15) with $\dot{\phi}_{B}>0$, $V_{0}=1.96\times 10^{-15}$ and $\alpha=0.0962$.
@@
Figure 20: The equation of state $w(t)$ with different values of $\phi_{B}$ for the the generalized T-Models given by (3.15) with $\dot{\phi}_{B}>0$, $V_{0}=1.96\times 10^{-15}$ and $\alpha=0.0962$.
@@
Figure 21: The numerical solutions of $a(t)$ and $\phi(t)$ for various initial conditions of $\phi_{B}$ for the generalized T-Models given by (3.15) with $\dot{\phi}_{B}>0$, $V_{0}=1.96*10^{-15}$ and $\alpha=0.0962$.